OpenStudy (anonymous):

A basket has 10 pencils, 6 pens, and 5 highlighters. Among those are 1 blue pencil, 1 blue pen, and 1 blue highlighter. You choose 1 of each item. What is the probability that the pencil and highlighter you pick are blue?

OpenStudy (anonymous):

@mathmate @jim_thompson5910 @Denebel

OpenStudy (mathmate):

Probability of picking the blue pencil out of 10 pencils is 1/10. Probability of picking the blue highlighter out of 5 highlighters is 1/5. If both events are to occur, and assuming the two events are independent, we can apply the multiplication rule: P(Blue Pen, Blue Highlighter) = (1/10)*(1/5) = 1/50

OpenStudy (anonymous):

So would it be (1/10)*(1/5)*(5/6)? 5/6 because you still need to choose one of each?

OpenStudy (mathmate):

Almost. It would be (1/10)*(1/5)*(6/6)= still 1/50. 6/6 because whatever you choose for pens does not matter. I did not read the line that said one of each item, but it comes out the same anyway.

OpenStudy (anonymous):

Thank you!

OpenStudy (mathmate):

You're welcome! :)